# No License

### v3

@@ Appendix C Model run time
As an additional evaluation, the run time per epoch for each model in the synthetic datasets is presented in Tables 6 and 7. The results, when considered alongside Tables 1 and 2, illustrate that SIR-GCN achieves a balance between computational complexity and model expressivity. Notably, SIR-GCN is computationally less expensive than PNA while delivering superior performance. These results further underscore the practicality of the proposed model.
Table 6: Run time per epoch on DictionaryLookup.
@@
Table 7: Run time per epoch on GraphHeterophily.
@@
Appendix D SIR-GCN extensions
Denote $\bm{h_{u,v}}$ as the feature of the edge connecting node $v$ to node $u$. Following the intuition presented in Eqns. 8 and 9, SIR-GCN with residual connection may be modified to leverage edge features to obtain
$\bm{h_{u}^{*}}=\textsc{MLP}_{\text{Res}}(\bm{h_{u}})+\sum_{v\in\mathcal{N}(u)}\bm{W_{R}}~{}\sigma\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{E}}\bm{h_{u,v}}+\bm{W_{K}}\bm{h_{v}}\right),$ (31)
where $\bm{W_{E}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{in}}}$. Consequently, this also increases the computational complexity of the model to
@@ Appendix D SIR-GCN extensions
with $\texttt{MLP}_{\text{Res}}$ denoting the computational complexity of $\textsc{MLP}_{\text{Res}}$, which is comparable to PNA. Similarly, this extension may be viewed as a generalization of GIN with edge features (Hu et al., 2019).
Furthermore, one may inject inductive bias into the pseudometrics which may correspond to specifying the architecture type for the corresponding message function $g$. For instance, if node features are known to have a sequential relationship (e.g., stock (Hsu et al., 2021) and fMRI (Kim & Ye, 2020) data), $g$ may then be aptly modeled using recurrent or convolutional networks.

### v4

@@ Appendix D Runtime analysis
As an additional evaluation, the inference runtime for each model in the synthetic datasets is presented in Tables 6 and 7. The results, when considered alongside Tables 1 and 2, illustrate that SIR-GCN achieves a balance between computational complexity and model expressivity, specifically with regards to PNA which is also designed for uncountable node features.
Table 6: DictionaryLookup inference runtime.
@@
Table 7: HeteroEdgeCount inference runtime.
@@
Table 8 further complements these results by presenting the asymptotic computational runtime complexity of the different GNNs. In particular, SIR-GCN first computes the linear transformations $\boldsymbol{W_{Q}}\boldsymbol{h_{u}}$ and $\boldsymbol{W_{K}}\boldsymbol{h_{v}}$ for every node which incurs $\mathcal{O}\left(\left|\mathcal{V}\right|\times d_{\text{hidden}}\times d_{\text{in}}\right)$. Afterward, $\sigma\left(\boldsymbol{W_{Q}}\boldsymbol{h_{u}}+\boldsymbol{W_{K}}\boldsymbol{h_{v}}\right)$ is computed for every edge, using the previously calculated values, and aggregated across the neighbors of each node which incurs $\mathcal{O}\left(\left|\mathcal{E}\right|\times d_{\text{hidden}}\right)$. Finally, the aggregated values are linearly transformed with $\boldsymbol{W_{R}}$ for every node which incurs $\mathcal{O}\left(\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{hidden}}\right)$ [^4] In the case of SIR-GCN with max aggregation, the linear transformation and the max aggregator cannot be interchanged. Hence, the linear transformation $\boldsymbol{W_{R}}$ must be performed along edges which incurs $\mathcal{O}\left(\left|\mathcal{E}\right|\times d_{\text{out}}\times d_{\text{hidden}}\right)$.. In total, since SIR-GCN employs only linear transformations along nodes and only an activation function along edges, its computational complexity is comparable to GCN, GraphSAGE, GAT, GATv2, and GIN. Specifically, these models achieve computational efficiency by maintaining linear complexity along edges attributed to activation functions and neighborhood aggregation. Despite this, SIR-GCN consistently outperforms these classical GNNs across all benchmarks. Notably, SIR-GCN also demonstrates a lower complexity than PNA due to the number of aggregators used, yet delivers superior performance across most datasets. These additional analyses further underscore the practical utility of the proposed model.
Table 8: Asymptotic runtime complexity.
Model Complexity
GCN $\mathcal{O}\left(\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{in}}+\left|\mathcal{E}\right|\times d_{\text{out}}\right)$
GraphSAGE $\mathcal{O}\left(\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{in}}+\left|\mathcal{E}\right|\times d_{\text{out}}\right)$
GAT/GATv2 $\mathcal{O}\left(\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{in}}+\left|\mathcal{E}\right|\times d_{\text{out}}\right)$
GIN $\mathcal{O}\left(\left|\mathcal{E}\right|\times d_{\text{in}}+\left|\mathcal{V}\right|\times\texttt{MLP}\right)$
PNA $\mathcal{O}\left(\left|\mathcal{E}\right|\times d_{\text{in}}^{2}+\left|\mathcal{E}\right|\times d_{\text{in}}\times k+\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{in}}\times k\right)$
SIR-GCN $\mathcal{O}\left(\left|\mathcal{V}\right|\times d_{\text{hidden}}\times d_{\text{in}}+\left|\mathcal{E}\right|\times d_{\text{hidden}}+\left|\mathcal{V}\right|\times d_{\text{out}}\times d_{\text{hidden}}\right)$
Note: $k$ represents the number of aggregators and scalers in PNA.
Appendix E SIR-GCN extensions
Denote $\boldsymbol{h_{u,v}}$ as the feature of the edge connecting node $v$ to node $u$. Following the intuition presented in Section 3.1, SIR-GCN with residual connection may be modified to leverage edge features to obtain
$\boldsymbol{h_{u}^{*}}=\textsc{MLP}_{\text{Res}}(\boldsymbol{h_{u}})+\sum_{v\in\mathcal{N}(u)}\boldsymbol{W_{R}}~{}\sigma\left(\boldsymbol{W_{Q}}\boldsymbol{h_{u}}+\boldsymbol{W_{E}}\boldsymbol{h_{u,v}}+\boldsymbol{W_{K}}\boldsymbol{h_{v}}\right),$ (31)
where $\boldsymbol{W_{E}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{in}}}$. Consequently, this also increases the computational complexity of the model to
@@ Appendix E SIR-GCN extensions
where $\texttt{MLP}_{\text{Res}}$ denotes the computational complexity of $\textsc{MLP}_{\text{Res}}$, making it comparable to PNA. Similarly, this extension may be viewed as a generalization of GIN with edge features (Hu et al., 2020b).
Furthermore, one may also inject inductive bias into the pseudometrics which may correspond to specifying the architecture type for the corresponding message function $g$. For instance, if node features are known to have a sequential relationship (e.g., stock (Hsu et al., 2023) and fMRI (Kim & Ye, 2020) data), $g$ may then be aptly modeled using recurrent or convolutional networks.